OpenStudy (anonymous):

Write the equation in point-slope form for a line that has a slope of 4 and passes through the point (3, 5).

OpenStudy (anonymous):

Alright, so we have the following as given: - m = 4 - x=3, y = 5 plug these in the following equation: y-yo= m(x-xo) and compute ^_^ give it a try now :)

OpenStudy (anonymous):

okay, thaaankyouu!(: